# Quantum entanglement in phase space

### v1

@@ Abstract
Testing criterion (III) can be performed in two ways. The first is to measure it in its original form, which is equivalent to performing single-mode tomography to obtain a full slice of the two-mode Wigner function. Alternatively, if the transformation $(x^{\prime},p^{\prime})$ corresponds to a unitary transformations acting on $\rho_{B}$, a second, potentially weaker method can be applied, requiring to measure only one point of the reduced Wigner function. Specifically, let us consider a beam-splitter transformation of the modes, $\hat{a}_{\pm}=(\hat{a}_{A}\pm\hat{a}_{B})/\sqrt{2}$. Using Eq. (11) with $\theta=\pi/4$ and $\left(x^{\prime}=\sqrt{2}x_{+}-x,p^{\prime}=\sqrt{2}p_{+}-p\right)$, we arrive at the special case $\iint_{-\infty}^{\infty}W_{AB}\left(x,p,x^{\prime},p^{\prime}\right)\text{d}x\text{d}p=\frac{1}{2}W_{+}\left(x_{+},p_{+}\right)$, which is half the Wigner function of one of the two modes resulting from the beam-splitter transformation. This means that, to violate criterion (III) it is enough to find a single point in the $(x_{+},p_{+})$ space where $W_{+}$ is negative [10, 11], see SM Sec. IV [22].
Applications and comparison to known criteria.– In the following, we illustrate how criteria (I,II,III) are applied to several experimentally relevant scenarios. For a comparison between our entanglement criteria and nonlocality/steering criteria based on the Wigner function see SM Sec.VI [22].
In the simplest case of TMSV states, i.e. $\eta=1,r=0$, we find that the linear transformation $(x^{\prime}=x,p^{\prime}=-p)$ and $\theta=\pi/4$ are optimal and allow us to detect entanglement with both criteria (I,II) as long as $s>0$ (See SM Sec.V [22]). This result coincides with the conclusion obtained from Duan-Simon and PPT criteria, which are known to be necessary and sufficient [1, 2, 20, 25].
In the most general case of a TMST state, i.e. $01>\eta>0,r>$ 0, Duan-Simon and PPT criteria are still necessary and sufficient conditions for entanglement [1, 2, 20, 25]. Remarkably, we can prove that this is also true for our criterion (I), see SM Sec.III [22]. On the other hand, while our criterion (II) is still useful, the bound it provides is not tight. In fact, for Gaussian states criterion (II) is a specific case of criterion (I) with $\theta=\pi/4$ and $R=\mathbb{R}^{2}$, and therefore (I) is strictly stronger than (II). In Fig. 2, we illustrate this with a concrete example. The shaded region indicates TMST states that are entangled. These states are all detected by Simon’s criterion, which is violated above the black solid line, since it is necessary and sufficient for Gaussian states [1]. The same conclusion can be obtained from our criterion (I). In addition, our criterion (II) detects only a fraction of all TMST entangled states, as indicated by the regions above the dashed coloured lines. Note that these depend on the squeezing $s$, while the bound obtained from criterion (I) is independent of $s$. Contrary to (I,II), entanglement criteria in Refs. [10, 11] fail to detect entanglement for any TMST states, more precisely, any Gaussian state, since they only work for states with Wigner negativity.
Werner states.– We consider here mixtures between a pure state |$\ket{\psi}$ and white noise, taking the form $\rho(\psi)=\epsilon|\psi\rangle\langle\psi|+(1-\epsilon)\mathbbm{1}/4$, where $\epsilon\in[0,1]$ specifies the purity. In the following, we will focus on two cases: $|\psi\rangle=|\Phi^{+}\rangle\equiv\left(|00\rangle+|11\rangle\right)/\sqrt{2}$ and $|\psi\rangle=|\Psi^{+}\rangle\equiv\left(|01\rangle+|10\rangle\right)/\sqrt{2}$.
Let us start with Werner state $\rho(\Phi^{+})$, whose Wigner distribution $W_{AB}$ under the linear transformation $(x^{\prime}=x,p^{\prime}=-p)$ is non-negative for any $\epsilon$ (See SM Sec.V [22]). Considering the integration to be over the entire space $\mathbb{R}^{2}$ and $\theta=\pi/4$, we obtain for criterion (II)
which is violated for $\epsilon>1/3$. The same conclusion is obtained from criterion (I). This result coincides with the known tight bound for entanglement of $\rho(\Phi^{+})$ [25].
For Werner state $\rho(\Psi^{+})$, we find that the Wigner function $W_{AB}(x,p,-x,-p)$ is negative over the entire phase space at $\epsilon=1$ (See SM Sec.V [22]). We thus calculate criterion (III) and obtain
Dephased two-mode cat states.
Figure 3: Entanglement detection for dephased cat states Eq. (16). (a): Slice of the joint Wigner function defined by $W_{AB}(x,p,x,-p)$, for $\gamma=2$ and $\epsilon=1$. Yellow dashed circles indicate the smallest region $R$ needed to violate criterion (II). (b): Amount of violation for criterion (II) with $\theta=\pi/4$, i.e. $\iint_{\infty}^{\infty}|W_{AB}(x,p,x,-p)|\text{d}x\text{d}p-1/(4\pi)$, and for PPT criterion, i.e. $|\lambda_{\min}(\rho^{T_{B}})|$, as a function of $\epsilon$ and for three values of $\gamma$. Entanglement can be revealed by both criteria whenever $\gamma,\epsilon>0$.
To illustrate the performance of our criteria on non-Gaussian states with many excitations, we consider states of the form
As illustrated in Fig. 3(a), the joint Wigner distribution under the transformation $(x^{\prime}=x,p^{\prime}=-p)$ and $\theta=\pi/4$, which we numerically find to be optimal, is positive over the entire space. When computing criteria (I,II) and PPT criterion we find that they are violated for any $\gamma,\epsilon>0$ (See SM Sec.V [22]). The value for the criteria is shown in Fig. 3(b). For the cat states considered, our criteria (I,II) are more powerful in detecting entanglement than the entanglement criterion discussed in Refs. [10, 11], since the latter has the same entanglement bound as criterion (III), which is useful but not tight for $\rho(\psi^{+})$ (See SM Sec.V [22]). Also in practical considerations, different from PPT criterion and entanglement criterion based on quantum Fisher information [26] where full information of the density matrix is required, the criteria (I,II) only need partial information (a slice of Wigner function) determined by the transformation $\left(x^{\prime},p^{\prime}\right)$. Additionally, for criterion (II) the required integral region can be further narrowed down to a limited region, for example, the yellow circles in Fig. 3(a).
Summary and outlook.– We have introduced criteria to detect CV bipartite entanglement from the Wigner function. By comparing them with known criteria from the literature [1, 2], we have shown cases in which they are necessary and sufficient. Crucially, our criteria require only partial information on the system’s state, and can be implemented experimentally by direct Wigner function measurements. Importantly, the amount of information required is no more than a single-mode Wigner tomography.
It would be interesting to investigate whether criteria similar to (I,II,III) can be found for discrete variable systems, by using the Wigner function for spin states [27]. This would allow for alternative methods for the detection of entanglement between e.g. atomic ensembles [28, 29, 30, 31, 32].
@@
Acknowledgments.– We are indebted to M. Walschaers for the inspiring and insightful discussions on the topic. This work is supported by the National Natural Science Foundation of China (Grants No. 12125402, No. 12405005), Beijing Natural Science Foundation (Grant No. Z240007), and the Innovation Program for Quantum Science and Technology (No. 2021ZD0301500). S.L. acknowledges the China Postdoctoral Science Foundation (No. 2023M740119). J. Guo acknowledges Postdoctoral Fellowship Program of CPSF (GZB20240027). M.F. was supported by the Swiss National Science Foundation Ambizione Grant No. 208886, and by The Branco Weiss Fellowship – Society in Science, administered by the ETH Zürich.
@@
I I. An intuitive phase-space counterpart to the Duan-Simon criterion and their comparison
@@
Using Williamson’s theorem, the covariance matrix of a single-mode Gaussian state can be diagonalized as $\sigma=\operatorname{diag}(l,l)$ through a symplectic transformation, where $l\geqslant 1$ due to the uncertainty relation. Conversely, if $10<l<$ 1 the purity $1/\sqrt{\operatorname{Det}\sigma}$ would exceed $1$. Consequently, from the purity of the partially transposed state we can establish an entanglement criterion based on the phase space distribution.
@@
II II. Detailed proof of criterion (I)
@@
where in the last row we have applied the substitution $\left(x\rightarrow x+\frac{\sin\theta}{\cos\theta}X,p\rightarrow p+\frac{\sin\theta}{\cos\theta}P\right)$. Now, assume that $\rho_{AB}$ is a separable pure state, meaning $\mathbbm{1}\otimes h_{2}h_{1}(\rho_{AB})$ should also correspond a physical state. Next, perform a beam-splitter transformation given by $\left(\begin{array}[]{cc}\cos\theta&\sin\theta\\ \sin\theta&-\cos\theta\end{array}\right)$ on the two modes of $\mathbbm{1}\otimes h_{2}h_{1}(\rho_{AB})$. The resulting Wigner function is given by $W_{Ah_{2}h_{1}\left(B\right)}\left(\cos\theta x+\sin\theta X,\cos\theta p+\sin\theta P,\sin\theta x-\cos\theta X,\sin\theta p-\cos\theta P\right)$ where $A^{\prime}:\left(x,p\right)$ and $B^{\prime}:\left(X,P\right)$ represent the two output modes. Therefore, Eq. (27) becomes the Wigner function of a reduced mode, which is upper-bounded by $1/(2\pi)$,
@@
III III. Criterion (I) is sufficient and necessary for Gaussian states
@@
IV IV. A comparison between criterion (III) and previous criteria based on Wigner functions
@@
V V. Detailed analysis of some exemplary quantum states
V.1 A. two-mode squeezed thermal states
@@
Then let’s analyze the most general case, i.e. $01>\eta>0,r>$ 0, where Simon’s criterion is still a necessary and sufficient conditions for detecting entanglement [1]. We have proved in SM Sec. III that our criterion (I) is necessary and sufficient for Gaussian states, so criterion (I) provides a tight bound for detecting the entanglement for TMST states (See Fig. 2 in the main text).
@@
In conclusion, our criteria, which perform as well as PPT criterion, are necessary and sufficient conditions for the considered Werner states. In addition, we have analyzed some known entanglement criteria for better comparison in Fig. 6. It’s found that Duan-Simon criteria [1, 2] and Hillery-Zubairy criteria [33] are unable to detect entanglement for any $\epsilon$, and the entanglement based on quantum Fisher information [26] is useful but not tight (it reveals entanglement for both $\rho(\Phi^{+})$ and $\rho(\Psi^{+})$ when $\epsilon\gtrsim 0.65$).
V.3 C. partially dephased two-mode cat states
Figure 5: Wigner distributions and entanglement detection of partially dephased two-mode cat state $\rho(\psi^{-})$. (a) The Wigner function $W_{AB}(x,p,x,p)$ (left) and $W_{AB}(x,p,-x,-p)$ (right) at $\epsilon=1$. (b) The minimum parameter $\epsilon_{\min}$ required to violate the criteria (II,III) as a function of $\gamma$. Note $\gamma=0$ gives a singular point.

### v2

@@ Methods
Acknowledgments.– We are indebted to M. Walschaers for inspiring and insightful discussions on the topic and to L. H. Zaw and E. Agudelo for valuable comments on our work. This work is supported by the National Natural Science Foundation of China (Grants No. 12125402, No. 12405005, No. 12447157), the Beijing Natural Science Foundation (Grant No. Z240007), the Innovation Program for Quantum Science and Technology (No. 2021ZD0301500), and the China Postdoctoral Science Foundation (No. 2023M740119 and 2024M760072). J. Guo acknowledges the CPSF Postdoctoral Fellowship Program (No. GZB20240027). MF was supported by the Swiss National Science Foundation Ambizione Grant No. 208886, and by the Branco Weiss Fellowship – Society in Science, administered by the ETH Zürich.
@@
I I. Criterion (I) can lower bound the entanglement negativity
If we consider a beam-splitter transformation $S$ and a time reversal $(x,p)\rightarrow(x,-p)$, the left-hand side of Criterion (I) can be re-expressed as
$\iint_{-\infty}^{\infty}\text{d}x\text{d}pW(\cos{\theta}x,\cos{\theta}p,\sin{\theta}x^{\prime},\sin{\theta}p^{\prime})$ $=\iint_{-\infty}^{\infty}\text{d}x\text{d}pW(\cos{\theta}x,\cos{\theta}p,\sin{\theta}x,-\sin{\theta}p)$
$=\iint_{-\infty}^{\infty}\text{d}x\text{d}pW_{\rho^{T_{B}}}(\cos{\theta}x,\cos{\theta}p,\sin{\theta}x,\sin{\theta}p)$
$=\iint_{\infty}^{\infty}\text{d}x\text{d}pW_{S(\rho^{T}_{B})}(x,p,0,0)$
$=W_{\text{tr}_{A^{\prime}}[S(\rho^{T_{B}})]}(0,0).$ (18)
Such Wigner function can be expressed by the displaced parity operators $\hat{P}(\alpha)=\hat{D}^{\dagger}(\alpha)(-1)^{\hat{n}}\hat{D}(\alpha)$
$2\pi W_{\text{tr}_{A^{\prime}}[S(\rho^{T_{B}})]}(0,0)$ $=\text{tr}\left[\hat{P}_{B^{\prime}}(0)\text{tr}_{A^{\prime}}[S(\rho^{T_{B}})]\right]$
$=\text{tr}\left[(\hat{I}_{A^{\prime}}\otimes\hat{P}_{B^{\prime}}(0))\cdot S(\rho^{T_{B}})\right].$ (19)
Here, $(\hat{I}_{A^{\prime}}\otimes\hat{P}_{B^{\prime}})$ and the beam-splitter transformation S can be regarded as a unitary operator $\hat{U}^{\prime}$. We can therefore obtain
[…]
Thus, We conclude that Criterion (I) provides a lower bound of the entanglement negativity $\mathcal{N}$
[…]
$\geq\pi\iint_{-\infty}^{\infty}\text{d}x\text{d}pW(\cos{\theta}x,\cos{\theta}p,\sin{\theta}x,-\sin{\theta}p)-\frac{1}{2}$ (21)
In fact, the bound is tight for Schmidt-correlated states $\sigma=\sum_{nm}c_{nm}|n,n\rangle\langle m,m|$, where $c_{nm}$ are real coefficients. We will prove it in the following.
The partial trace on mode B yields $\sigma^{T_{B}}=\sum_{nm}c_{nm}|n,m\rangle\langle m,n|$. The norm of $\sigma^{T_{B}}$ is $\sigma^{T_{B}}||\equiv\sqrt{\left(\sigma^{T_{B}}\right)^{\dagger}\left(\sigma^{T_{B}}\right)}=\sqrt{\sum_{n,m}c^{2}_{nm}|m,n\rangle\langle m,n|}=\sum_{n,m}c_{nm}|m,n\rangle\langle m,n|$. By multiplying a swap operator $S_{\text{swap}}=\sum_{n,m}|m,n\rangle\langle n,m|$, we have $\sigma^{T_{B}}S_{\text{swap}}=\sum_{n^{\prime},m^{\prime},n,m}c_{nm}(|n,m\rangle\langle m,n|)$ $(|m^{\prime},n^{\prime}\rangle\langle n^{\prime},m^{\prime}|)=\sum_{n,m}c_{nm}|m,n\rangle\langle m,n|$. Thus it is found that $\sigma^{T_{B}}||=\sigma^{T_{B}}S_{\text{swap}}$.
Therefore, we can further obtain
[…]
$=\iint_{-\infty}^{\infty}\text{d}x_{1}\text{d}p_{1}\text{d}x_{2}\text{d}p_{2}W_{\sigma^{T_{B}}}(x_{1},p_{1},x_{2},p_{2})\cdot 4\pi\delta\left(x_{1}-x_{2}\right)\delta\left(p_{1}-p_{2}\right)$
$=4\pi\iint_{-\infty}^{\infty}\text{d}x\text{d}pW_{\sigma^{T_{B}}}\left(x,p,x,p\right)$
$=4\pi\iint_{-\infty}^{\infty}\text{d}x\text{d}pW_{\sigma}\left(x,p,x,-p\right)$
$=2\pi\iint\text{d}x\text{d}pW_{\sigma}(x/\sqrt{2},p/\sqrt{2},x/\sqrt{2},-p/\sqrt{2}).$ (22)
Here, we use the Wigner function of a swap operator $4\pi\delta\left(x_{1}-x_{2}\right)\delta\left(p_{1}-p_{2}\right)$. The last raw is the left-hand side of Criterion (I) under the transformation with a time-reversal $(x,p)\rightarrow(x,-p)$ and $\theta=\pi/4$.
II II. Choice of $\Delta=\pm 1$ in linear transformations
II.1 A. Criterion (I) requires $\Delta=-1$
Criterion (I) can detect entanglement if and only if $\Delta=-1$. Now we prove it by demonstrating that Criterion (I) will not be violated for any states if $\Delta=+1$.
In the case of $\Delta=+1$, the quadratures ($x^{\prime},p^{\prime}$) is obtained from ($x,p$) through a unitary transformation, i.e. $x^{\prime}=U^{\dagger}xU,p^{\prime}=U^{\dagger}pU$. Then according to Eq. (37) and Eq. (38) of Sec. IV in SM, Criterion (I) can be expressed as
$\iint_{-\infty}^{\infty}W_{AB}(\cos{\theta}x,\cos{\theta}p,\sin{\theta}x^{\prime},\sin{\theta}p^{\prime})\text{d}x\text{d}p$
$=\iint_{-\infty}^{\infty}W_{Ah_{2}h_{1}(B)}\left(\cos{\theta}x+\sin{\theta}X,\cos{\theta}p+\sin{\theta}P,\sin{\theta}x-\cos{\theta}X,\sin{\theta}p-\cos{\theta}P\right)\text{d}x\text{d}p$
$=W_{B^{\prime}}(X,P),$
where $(X,P)$ are constant coordinates, and $B^{\prime}$ is an output mode of the beam-splitter with input being $\rho^{\prime}_{AB}=\mathbbm{1}\otimes h_{2}h_{1}(\rho_{AB})$. For any initial states $\rho_{AB}$, the state $\rho^{\prime}_{AB}$ after the unitary transformation is
$\rho^{\prime}_{AB}=\left(\mathbbm{1}\otimes U\right)\rho_{AB}\left(\mathbbm{1}\otimes U\right)^{\dagger}=U^{{}^{\prime}}\rho_{AB}U^{{}^{\prime}\dagger},$ (23)
where $h_{1},h_{2}$ are state transformations in Eqs. (35,36) and $U,U^{\prime}$ are unitary transformation on mode B and modes AB, respectively. After the unitary transformation, $\rho^{\prime}_{AB}$ should be a physical state, so that $B^{\prime}$ is also physical. Therefore, the Wigner function of the reduced mode $B^{\prime}:(X,P)$ is always upper-bounded by $1/(2\pi)$, meaning Criterion (I) will not be violated.
II.2 B. Criterion (III) requires $\Delta=+1$
Criterion (III) can detect entanglement if and only if $\Delta=+1$. We prove that Criterion (III) holds for all states if $\Delta=-1$.
$\Delta=-1$ indicates a mirror reflection, as an example we consider the time reversal $(x,p)\rightarrow(x,-p)$. The local linear coordinate transformation corresponds to partial transpose of the state $W_{\rho}(x_{A},p_{A},x_{B},-p_{B})=W_{\rho^{T_{B}}}(x_{A},p_{A},x_{B},p_{B})$. Based on Eq. (44), we obtain
$W_{\operatorname{tr}_{-}(\rho^{T_{B}})}\left(x_{+},p_{+}\right)=$ $\iint_{-\infty}^{\infty}W_{\rho^{T_{B}}}\left(\frac{x_{+}+x_{-}}{\sqrt{2}},\frac{p_{+}+p_{-}}{\sqrt{2}},\frac{x_{+}-x_{-}}{\sqrt{2}},\frac{p_{+}-p_{-}}{\sqrt{2}}\right)\text{d}x_{-}\text{d}p_{-}$
$=$ $2\iint_{-\infty}^{\infty}W_{\rho^{T_{B}}}\left(x_{-},p_{-},-x_{-}+\sqrt{2}x_{+},-p_{-}+\sqrt{2}p_{+}\right)\text{d}x_{-}\text{d}p_{-}$
$=$ $2\iint_{-\infty}^{\infty}W_{\rho^{T_{B}}}\left(x,p,-x+x_{0},-p+p_{0}\right)\text{d}x\text{d}p$
$=$ $2\iint_{-\infty}^{\infty}W_{\rho}\left(x,p,-x+x_{0},p-p_{0}\right)\text{d}x\text{d}p$
$=$ $2\iint_{-\infty}^{\infty}W_{\mathbbm{1}\otimes f(\rho)}\left(x,p,x,-p\right)\text{d}x\text{d}p,$ (24)
where $f:(x,-p)\rightarrow(-x+x_{0},p-p_{0})$ corresponds to a local unitary.
In [12], the author pointed out that the negativity of the reduced Wigner function implies partial transpose negativity. For $\rho^{T_{B}}$, Theorem 2 in [12] gives
$W_{\operatorname{tr}_{-}(\rho^{T_{B}})}(x_{+},p_{+})\ngeq 0\Rightarrow\left(\rho^{T_{B}}\right)^{T_{B}}=\rho\nsucceq 0,$ (25)
which obviously conflicts with the fact that for any physical state $\rho\succeq 0$. Therefore, we obtain the following conclusion that
$W_{\operatorname{tr}_{-}(\rho^{T_{B}})}(x_{+},p_{+})\geq 0,$ (26)
which means Criterion (III) holds for any states.
@@
III III. An intuitive phase-space counterpart to the Duan-Simon criterion and their comparison
@@
Using Williamson’s theorem, the covariance matrix of a single-mode Gaussian state can be diagonalized as $\sigma=\operatorname{diag}(l,l)$ through a symplectic transformation, where $l\geqslant 1$ due to the uncertainty relation. Conversely, if $0<l<1$ the purity $1/\sqrt{\operatorname{Det}\sigma}$ would exceed $1$. Consequently, from the purity of the partially transposed state we can establish an entanglement criterion based on the phase space distribution.
@@
IV IV. Detailed proof of criterion (I)
@@
where in the last row we have applied the substitution $\left(x\rightarrow x+\frac{\sin\theta}{\cos\theta}X,p\rightarrow p+\frac{\sin\theta}{\cos\theta}P\right)$. Now, assume that $\rho_{AB}$ is a separable pure state, meaning $\mathbbm{1}\otimes h_{2}h_{1}(\rho_{AB})$ should also correspond a physical state. Next, perform a beam-splitter transformation given by $\left(\begin{array}[]{cc}\cos\theta&\sin\theta\\ \sin\theta&-\cos\theta\end{array}\right)$ on the two modes of $\mathbbm{1}\otimes h_{2}h_{1}(\rho_{AB})$. The resulting Wigner function is given by
$W_{Ah_{2}h_{1}\left(B\right)}\left(\cos\theta x+\sin\theta X,\cos\theta p+\sin\theta P,\sin\theta x-\cos\theta X,\sin\theta p-\cos\theta P\right)$ where $A^{\prime}:\left(x,p\right)$ and $B^{\prime}:\left(X,P\right)$ represent the two output modes. Therefore, Eq. (37) becomes the Wigner function of a reduced mode, which is upper-bounded by $1/(2\pi)$,
@@
V V. Criterion (I) is sufficient and necessary for Gaussian states
@@
VI VI. A comparison between criterion (III) and previous criteria based on Wigner functions
@@
VII VII. Comparison with Bell inequality and EPR steering criterion based on Wigner functions
VII.1 A. EPR criterion based on pseudospin measurements
$W_{\Pi^{x}}(x,p)$ $=\text{sgn}(x),$
$W_{\Pi^{x}}(x,p)$ $=-\delta(x)\mathcal{P}\frac{1}{p},$ (46)
$W_{\Pi^{z}}(x,p)$ $=-\pi\delta(x)\delta(p),$ (47)
with $\mathbf{\xi}=(x_{A},p_{A},x_{B},p_{B})^{T}$.
VII.2 B. CHSH-type Bell inequality based on Wigner fucntions
VIII VIII. Detailed analysis of some exemplary quantum states
VIII.1 A. two-mode squeezed thermal states
@@
Then let’s analyze the most general case, i.e. $1>\eta>0,r>0$, where Simon’s criterion is still a necessary and sufficient conditions for detecting entanglement [1]. We have proved in SM Sec. V that our criterion (I) is necessary and sufficient for Gaussian states, so criterion (I) provides a tight bound for detecting the entanglement for TMST states (See Fig. 2 in the main text).
@@
In conclusion, our criteria, which perform as well as PPT criterion, are necessary and sufficient conditions for the considered Werner states. In addition, we have analyzed some known entanglement criteria for better comparison in Fig. 8. It’s found that Duan-Simon criteria [1, 2] are unable to detect entanglement for any $\epsilon$, and the entanglement based on quantum Fisher information [40] and Hillery-Zubairy criteria [50] are useful but not tight (the former reveals entanglement for both $\rho(\Phi^{+})$ and $\rho(\Psi^{+})$ when $\epsilon\gtrsim 0.65$, and the latter can only detect entanglement for $\rho(\Psi^{+})$ when $\epsilon>\frac{\sqrt{5}-1}{2}$ ).
VIII.3 C. partially dephased two-mode cat states
Figure 7: Wigner distributions and entanglement detection of partially dephased two-mode cat state $\rho(\psi^{-})$. (a) The Wigner function $W_{AB}(x,p,x,p)$ (left) and $W_{AB}(x,p,-x,-p)$ (right) at $\gamma=1,\epsilon=1$. (b) The minimum parameter $\epsilon_{\min}$ required to violate the criteria (II,III) as a function of $\gamma$. Note $\gamma=0$ gives a singular point.